Fermi National Laboratory


Accelerator Update

Wednesday August 4
The day shift began with Operations monitoring the Recycler stash, with Tevatron (TeV) experts conducting studies, with high voltage technicians working on the TeV sector E2 feeder, with an expert repairing an Antiproton Source (Pbar) Debuncher kicker, and with MiniBooNE and the Meson MTest experiment taking beam.

Operations reported at 9:16 AM that a tech had completed his repairs on the Debuncher kicker (D:EKIK).

Pbar began stacking at 9:45 AM.

Operations reported at 10:12 AM that they had halted stacking so experts could check something on EKIK.

Operations reported that Controls personnel had rebooted the Time Line Generator (TLG) at 10:33 AM.

Pbar resumed stacking at 12:17 PM. Operators reported that EKIK could be turned off remotely, but not on.

Operations put the TeV into shot setup at 1:58 PM.

At 2:47 PM, an Accumulator power supply (A:QDF) tripped off and the stack was lost. Later it was discovered that a person had opened a power supply door, which caused the supply to trip off.

Operations made Pbar ready for access at 3:11 PM. They came out at 4:53 PM.

Operations tried to put Pbar into stacking mode at 5:10 PM, but had trouble closing a Debuncher contactor (D:IB). Operators called an expert who started work at 6:14 PM and guessed at a two-hour repair. He actually had it finished by 7:25 PM.

Operations reported at 8:14 PM that they had beam to Pbar.

Thursday August 5
The midnight shift began with TeV expert conducting studies, with Pbar stacking, and with MiniBooNE taking beam.

MTest experimenters requested beam at 12:52 AM and Operations sent it.

TeV studies ended at 4:44 AM.

Operations put the TeV through a dry squeeze at 5:03 AM.

Operations put the TeV into shot setup at 5:15 AM. This disabled all Switchyard (SY) beam.

A Booster RF station (BRF13) tripped off at 5:15 AM for the first of many times in twenty minutes. Operators bypassed the station.

Operations prepared MiniBooNE for an access at 7:35 PM.

Since there is no stacking during shot setup, an expert continued working on EKIK at 7:38 AM.

At 8 AM, the TeV shot setup was put on hold to make repairs to a TeV RF station (TRF3). This work ended at 10:33 AM and shot setup resumed about two minutes later.

Operations reported that Linac was ready for access at 11:46 AM.

Operations had Booster ready for access by 11:52 AM.

Operations established store 3711 at 11:57 AM with an initial luminosity of 45.21. This was a Mixed Antiproton Mode shot. The low luminosity seemed to be due to an emittance problem with the first transfer.

Operations had the Main Injector (MI) ready for access at 1:30 PM.

Linac came out of access at 2:22 PM and an expert began tuning the beam line thirty minutes later.

MI came out of access at 3:06 PM.

Operations reported at 7:14 PM that RF techs were still working on two Booster RF stations. These stations were bypassed. Operations began the Booster turn on.

Operations established beam through Booster and to the Pbar Debuncher at 8:39 PM, but D:EKIK wouldn't fire. Operators investigated and found that the kicker's ramp tables were corrupted. Operator loaded new tables and Pbar resumed stacking around 9 PM.

Operations established beam to MiniBooNE at 9:22 PM.

Operations established beam to the Meson MTest experiment at 9:57 PM.

Operations reported at 10:50 PM that the RF techs had finished the work on BRF8 & 13. Operators tied them back into the Booster RF system.

Operations established beam to MCenter at 11:45 PM.

Friday August 6
The midnight shift began with Operations monitoring store 3711, with Pbar stacking, and with MiniBooNE, MCenter, and MTest taking beam.

The cryo system techs at CHL reported at 1:53 AM that they have been fighting to keep the TeV sector D1 cryo system operating. Operators warned the experiments that they might have to abort the beam before a quench occurred. The Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments took precautions to protect their detectors.

Operations reported at 3:31 AM that the D1 cryo system difficulties might be caused by a frozen vacuum o-ring. CHL said that they had stabilized the system.

CDF and D0 reported at 3:41 AM that they had resumed taking data.

At 4:44 AM, Operations received alarms from the idle Neutrino beam lines in the Fixed Target area. (Meson is the only Fixed Target area that is able to receive beam.) They also received a low-pressure alarm from CUB.

Operations reported at 5:48 AM that feeder 32 faulted. This is what caused the Neutrino and CUB alarms. The High Voltage Coordinator was called in to investigate.

The High Voltage Coordinator reported at 7:11 that he had transferred the power load from feeder 32 to feeders 33 and 35.

Plans for Friday and the Weekend
The plan for today is to terminate the beam at 9 AM. TeV vacuum techs will access sector D1 and investigate the vacuum problem. If the repair goes quick, the Run Coordinator will allow TeV expert to conduct some studies before the next store. The plan for the weekend is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/6/2004   email Fermilab